A kitchen sink is filled with water. A dripping kitchen faucet creates circular water waves along the water’s surface which travel at a speed of 0.042 m/s and have a wavelength of 0.05 m. How many drips strike the water’s surface in 12 s?


1
Expert's answer
2021-01-27T09:30:55-0500

"\lambda=\frac{v}{\nu}\to \ \nu=\frac{v}{\lambda}=\frac{0.042}{0.05}=0.84(Hz)"


"n=0.84\cdot12\approx10" . Answer